Information pertaining to Luigi Co. is as follows:

Sales
Variable cost of goods sold Direct fixed overhead

Fleece Blanket

$425,000 124,000 29,000

Thermal Blanket

$524,000 213,000 36,000

A sales commission of 3% of sales is paid for each of the two product lines. Direct fixed selling and administrative expense was estimated to be $32,000 for the fleece blankets and $66,000 for the thermal blankets. Common fixed overhead for the factory was estimated to be $83,000 and common selling and administrative expense was estimated to be $14,000.

Required: Prepare a segmented income statement for Luigi Co. for the coming year, using variable costing.
Answer #1

Answer-

LUIGI CO.
CONTRIBUTION FORMAT SEGMENTED INCOME STATEMENT
USING VARIABLE COSTING
PARTICULARS FLEECE BLANKET THERMAL BLANKET TOTAL COMPANY
$ $ $
Sales(a) 425000 524000 949000
Less - Variable expenses
Cost of goods sold 124000 213000 337000
Sales commissions @ 3% 12750 15720 28470
Total variable expenses (b) 136750 228720 365470
Contribution margin (c=a-b) 288250 295280 583530
Traceable fixed expenses
Direct overhead 29000 36000 65000
Selling & administrative expenses 32000 66000 98000
Total traceable fixed expenses (d) 61000 102000 163000
Product line segment margin (e=c-d) 227250 193280 420530
Less- Common fixed expenses
Direct overhead 83000
Selling & administrative expenses 14000
Total common fixed expenses (f) 97000
Net income (g= e-f) 323530